The sum of digit of a 2digit number is 9 the number obtained by interchanging the digits exceed the given number by 27 find the given number
Answer by nihar*2013(34)

First Number at unit place is is x and tens place is y
So original number is 10y+x
After interchnaging digits number will be 10x+y
number obtained by interchanging the digits exceed the given number by 27

10y+x+27 = 10x+y

10y-y+x-10+27=0
9y-9x+27=0
9 (3y-3x+9=0)
3y-3x+9=0---------(1)
The sum of digit of a 2digit number is 9
x+y= 9---------------(2)
can be solved further by any method i will do by elimination method
Multiplyinjg eq 1 by 1 3y-3x=9
Multiplyinjg eq 2 by 3 3y+3x=27
- - -
- 6x = -18
x= -18/-6
x= 3
Subtituting value in eq 1
x+y=9
3+y=9
y=6
If u need it in word format let me know i have Many sums like this solved in word 2010 that looks best thanks